OpenStudy (anonymous):

Medal and fan y^4*y^6= n^12/n^4= (b^3)^2

OpenStudy (anonymous):

law of indices states add the powers is base are the same

OpenStudy (anonymous):

so in y^4*y^6 =y^10

OpenStudy (anonymous):

ok and 2 would be n^15?

OpenStudy (anonymous):

for second one coz it division you subtract the power n^12/n^4=n^8

OpenStudy (anonymous):

multiply the power in the last question =b^6

OpenStudy (anonymous):

ok think I understand

OpenStudy (anonymous):

three rules when multiplying add the power division, subtract the powers and in bracket multiply the powers